OpenStudy (liv33):

Solve the equation: 6 = x+2/3

satellite73 (satellite73):

\[6=\frac{x+2}{3}\] right?

OpenStudy (liv33):

yes

satellite73 (satellite73):

a) multiply both sides by 3 to get rid of the fraction, then b) subtract 2 from both sides

OpenStudy (liv33):

How would I multiply 3 on the fraction side?

satellite73 (satellite73):

cancel the 3, that is the whole point of multiplying by three, gets rid of the three in the denominator

satellite73 (satellite73):

\[\frac{x+2}{3}\times 3=\frac{x+3}{\cancel3}\times \cancel 3=x+2\]

OpenStudy (liv33):

Oh ok, I see now

OpenStudy (liv33):

It then becomes 18 = x + 2 ?

satellite73 (satellite73):

yes

OpenStudy (liv33):

Then 16 = x

satellite73 (satellite73):

yes

OpenStudy (liv33):

Ahh ok

OpenStudy (liv33):

Thanks :)

satellite73 (satellite73):

yw